# Enhanced Monocular Visual Odometry with AR Poses and Integrated INS-GPS for Robust Localization in Urban Environments

### v1

@@ III-F1 Visual Odometry and AR Poses
Visual odometry is estimating the motion of the vehicle using visual cues from the cameras attached to it. We deployed a featured based monocular odometry. Our pipeline for visual odometry is as shown in Fig. 5. We use the FAST corner detection algorithm [14][15] to detect features in each frame It due to it’s fast computational efficiency. Each of these corner pixels are tracked using the optical flow bases KLT tracker [16] in the successive frame It+1. By selecting points using RANSAC from pairs of corresponding points in successive frames we estimate the essential matrix E as shown in Eqn. 4 using the Nister five-point algorithm[17]. Furthermore, decomposing the essential matrix using SVD and exploiting rotation matrix constraints gives the rotation and translation matrix between the two consecutive frames as shown in Eqn. 5 and 6.
@@
Translation in Monocular visual odometry suffers from scale factor issues that need to be estimated using external sensors. We use AR poses as obtained from the Augmented Reality Toolkit present in iPhone. The relative scale is evaluated by calculating the Frobenius norm between the transnational vectors of the two frames.
@@ III-F2 INS and Estimated GPS
Initial Navigation system (INS) is a state estimation system that uses data from a IMU to continuously calculate the attitude of the system by dead reckoning (Position, Velocity and Orientation). IMU consists of a accelerometer, gyroscope and magnetometer that measures the linear acceleration, angular velocity and magnetic field of the Vehicle respectively for all 3 directions. We used a strapdown INS that integrates angular rate in attitude estimation and uses the attitude data to transform the IMU acceleration into latitude, longitude and heading. A detailed explanation and equations used for INS are given here [19][22]. The output trajectory is prone to gyroscope and accelerometer bias and random noise that gets added during the dead reckoning process to cause a drift from the actual trajectory. We used an Extended Kalman Filter that inputs high-frequency IMU data and predicts the state of the system based on the INS one time stamp ahead. It also updates the error covariance of the EKF filter. For each low-frequency GPS data we update filter state and error covariance by computing the Kalman gain. The state for the system is given as: $x=\begin{bmatrix}q_{0}&q_{1}&q_{2}&q_{3}&pos_{N}&pos_{E}&V_{N}&V_{E}\end{bmatrix}^{T}$ And the state space equation as:
@@ IV Testing Methodology
Since we aim to show that our system performs well with the vast Google Street View dataset, we cannot test with standard autonomous driving datasets such as KITTI. This complicates direct comparisons to alternate methods and requires us to source our own ground truth. A common source of ground truth for global localization tasks is the use of an RTK-GPS system which are capable of centimeter-level accuracy in ideal conditions. However, we were not able to acquire an RTK system and instead rely on a manually annotated trajectory on top of a satellite map. For our data collection, we drove in a single lane with a front-facing camera and later approximated the camera trajectory with a piecewise linear segments. While it is difficult to ascertain the exact accuracy of this method, we expect the ground-truth trajectory to be within 1m of the true trajectory at any given point due to the width of the lane and position of the camera.
In order to evaluate our method of localization, we compare each of the estimated global locations represented by a latitude and longitude using the WGS-84 projection to our ground truth trajectory. We find the minimum distance from that point to the piecewise linear trajectory and then calculate the RMSE of a sequence of estimated points. Our estimated trajectory is formulated as a sequence of points, $T_{est}=\{(x_{0},y_{0}),\dots,(x_{n},y_{n})\}$ and our ground truth trajectory is similarly: $T_{truth}=\{(x_{0},y_{0}),\dots,(x_{m},y_{m})\}$.
We define our error function for any $p\in T_{est}$ in (8)
@@
where $0\leq i<j\leq m$, $i+1=j$, and $\operatorname{dist}(i,j,p))$ represents the minimum distance from $p$ to the line defined by the points $i,j$. We finally define the RMSE in (9):
@@
This method does not capture time information and thus would not tell us whether our estimated points are ahead or behind the ground truth trajectory in time. Similarly, since we do not have time information for our ground truth, we cannot plot the error over time and instead choose to plot over frames. This means our RMSE is dependent on factors such as vehicle speed, unexpected stops, etc. A possible approach might be to use the coarse GPS data to create a plot of error versus position but we chose to not implement this since it adds additional sources of error into our measurement that we cannot account.
Another unexpected source of error came from the panorama poses themselves, as provided by the Google Street View API. Despite the high-accuracy expected given the laser-range data and RTK-GPS used by Google Street View cars [7], we found significant deviations in some locations by comparing the reported coordinates to known locations on a map. While several satellite imagery sources showed similar deviations, it is difficult to determine whether the reported coordinates are erroneous or the map coordinates themselves but regardless, this error is incorporated when we measure the estimated trajectory deviation from ground truth.
@@ V Results
We chose to implement our pose graph optimization in both g2o and Ceres. We compare their performance given our setup in I over a varying subsets of matched features. As previously described, this value determines how many queried panoramas a given feature must be present in to be considered for bundle adjustment. We find that the g2o solver outperforms Ceres for all of the tested feature subsets, although Ceres appears to be much more robust to the varying features, producing similar results over the different subsets. For our fixed nonlinear BA, we find that taking the intersection of all 4 panoramas produces the best results.
Our primary metric as previously described is $RMSE(T_{est})$. We tested over a 1km track in Westwood, CA that saw variable traffic density, vehicle speed, and building density and distance. Over the entire track, we achieved an $RMSE(T_{est})=1.529m$ using our fixed pose bundle adjustment.
Looking at a plot of $\operatorname{err}(p)$ in Figure 10 we see significant variability in the results that appear to be due to the varying environmental conditions. We see frames 5800-6800 have a low error variance which corresponds to a period in which the vehicle is stopped at a light.
@@
A plot of the trajectory generated by IMU dead reckoning, panoramic GPS and INS + GPS integration is shown in Figure 10. From the plot we can see a significant bias in the gyroscope and accelerometer sensor that causes a drift in the trajectory. The INS + GPS trajectory near the panoramic GPS trajectory but at instances show vast deviation from the road when plotted and analysed visually. There is a scope of improvement in the trajectory generation by filtering out erroneous IMU data, initial calibration of the INS system and better filter tuning based on sensor characteristics.
@@ VI Future Work
There are several aspects of our system that we believe may be significantly improved upon. One area which we did not focus on in this paper is implementing our pipeline directly on a mobile phone running in real-time, however many of the components of our pipeline are already present in many mobile devices. Visual-Inertial Odometry and Extended Kalman Filters are, for example, present on most iOS devices with ARKit and run in real-time. Feature extraction and matching has been demonstrated to run in real-time at 30fps on mobile phones [13] and newer implementations of SIFT and nearest-neighbor matching on the GPU has shown these both of these steps can be performed in less than 4ms combined [10]. Performing real-time local bundle adjustment is more challenging, however although implementations have been demonstrated to run in 500ms on a mobile phone [12] and 230ms on a GPU [11]. Finally, while we found VLDs to perform robust outlier rejection for our pipeline, we are unaware of any work to accelerate K-VLD filtering or perform it on a mobile device. Future work may examine the performance of K-VLD filtering on mobile devices and the potential to perform these computations on a remote server and transmit the final results to the client mobile device.
@@ VII Conclusion
Our initial motivation for this project was that the sensor suite typically required for autonomous navigation is expensive and inaccessible for the vast majority of the population with motor vehicles. To address this issue, we built a system for accurate localization that is not gated by cost and would be easily accessible to anyone with a car. Using only a monocular camera, IMU, and GPS from a smartphone, along with panoramas retrieved from Google Street View, our method is able to compute a localization and trajectories with visual inertial odometry. Importantly, our method is able to localize with lane level accuracy, which is better than GPS alone, especially in dense city environments with large buildings or structures that obstruct RF signals. Additionally, with advancement in smartphone sensor technology and general smartphone availability, our method will only become more accurate and accessible.

### v2

@@ III-F1 Visual Odometry and AR Poses
Visual odometry is estimating the motion of the vehicle using visual cues from the cameras attached to it. I deployed a featured based monocular odometry. my pipeline for visual odometry is as shown in Fig. 5. I use the FAST corner detection algorithm [14][15] to detect features in each frame It due to it’s fast computational efficiency. Each of these corner pixels are tracked using the optical flow bases KLT tracker [16] in the successive frame It+1. By selecting points using RANSAC from pairs of corresponding points in successive frames I estimate the essential matrix E as shown in Eqn. 4 using the Nister five-point algorithm[17]. Furthermore, decomposing the essential matrix using SVD and exploiting rotation matrix constraints gives the rotation and translation matrix between the two consecutive frames as shown in Eqn. 5 and 6.
@@
Translation in Monocular visual odometry suffers from scale factor issues that need to be estimated using external sensors. I use AR poses as obtained from the Augmented Reality Toolkit present in iPhone. The relative scale is evaluated by calculating the Frobenius norm between the transnational vectors of the two frames.
@@ III-F2 INS and Estimated GPS
Initial Navigation system (INS) is a state estimation system that uses data from a IMU to continuously calculate the attitude of the system by dead reckoning (Position, Velocity and Orientation). IMU consists of a accelerometer, gyroscope and magnetometer that measures the linear acceleration, angular velocity and magnetic field of the Vehicle respectively for all 3 directions. I used a strapdown INS that integrates angular rate in attitude estimation and uses the attitude data to transform the IMU acceleration into latitude, longitude and heading. A detailed explanation and equations used for INS are given here [19][22]. The output trajectory is prone to gyroscope and accelerometer bias and random noise that gets added during the dead reckoning process to cause a drift from the actual trajectory. I used an Extended Kalman Filter that inputs high-frequency IMU data and predicts the state of the system based on the INS one time stamp ahead. It also updates the error covariance of the EKF filter. For each low-frequency GPS data I update filter state and error covariance by computing the Kalman gain. The state for the system is given as: $x=\begin{bmatrix}q_{0}&q_{1}&q_{2}&q_{3}&pos_{N}&pos_{E}&V_{N}&V_{E}\end{bmatrix}^{T}$ And the state space equation as:
@@ IV Testing Methodology
Since I aim to show that my system performs well with the vast Google Street View dataset, I cannot test with standard autonomous driving datasets such as KITTI. This complicates direct comparisons to alternate methods and requires me to source my own ground truth. A common source of ground truth for global localization tasks is the use of an RTK-GPS system which are capable of centimeter-level accuracy in ideal conditions. However, I were not able to acquire an RTK system and instead rely on a manually annotated trajectory on top of a satellite map. For my data collection, I drove in a single lane with a front-facing camera and later approximated the camera trajectory with a piecewise linear segments. While it is difficult to ascertain the exact accuracy of this method, I expect the ground-truth trajectory to be within 1m of the true trajectory at any given point due to the width of the lane and position of the camera.
In order to evaluate my method of localization, I compare each of the estimated global locations represented by a latitude and longitude using the WGS-84 projection to my ground truth trajectory. I find the minimum distance from that point to the piecewise linear trajectory and then calculate the RMSE of a sequence of estimated points. my estimated trajectory is formulated as a sequence of points, $T_{est}=\{(x_{0},y_{0}),\dots,(x_{n},y_{n})\}$ and my ground truth trajectory is similarly: $T_{truth}=\{(x_{0},y_{0}),\dots,(x_{m},y_{m})\}$.
I define my error function for any $p\in T_{est}$ in (8)
@@
where $0\leq i<j\leq m$, $i+1=j$, and $\operatorname{dist}(i,j,p))$ represents the minimum distance from $p$ to the line defined by the points $i,j$. I finally define the RMSE in (9):
@@
This method does not capture time information and thus would not tell me whether my estimated points are ahead or behind the ground truth trajectory in time. Similarly, since I do not have time information for my ground truth, I cannot plot the error over time and instead choose to plot over frames. This means my RMSE is dependent on factors such as vehicle speed, unexpected stops, etc. A possible approach might be to use the coarse GPS data to create a plot of error versus position but I chose to not implement this since it adds additional sources of error into my measurement that I cannot account.
Another unexpected source of error came from the panorama poses themselves, as provided by the Google Street View API. Despite the high-accuracy expected given the laser-range data and RTK-GPS used by Google Street View cars [7], I found significant deviations in some locations by comparing the reported coordinates to known locations on a map. While several satellite imagery sources showed similar deviations, it is difficult to determine whether the reported coordinates are erroneous or the map coordinates themselves but regardless, this error is incorporated when I measure the estimated trajectory deviation from ground truth.
@@ V Results
I chose to implement my pose graph optimization in both g2o and Ceres. I compare their performance given my setup in I over a varying subsets of matched features. As previously described, this value determines how many queried panoramas a given feature must be present in to be considered for bundle adjustment. I find that the g2o solver outperforms Ceres for all of the tested feature subsets, although Ceres appears to be much more robust to the varying features, producing similar results over the different subsets. For my fixed nonlinear BA, I find that taking the intersection of all 4 panoramas produces the best results.
my primary metric as previously described is $RMSE(T_{est})$. I tested over a 1km track in Westwood, CA that saw variable traffic density, vehicle speed, and building density and distance. Over the entire track, I achieved an $RMSE(T_{est})=1.529m$ using my fixed pose bundle adjustment.
Looking at a plot of $\operatorname{err}(p)$ in Figure 10 I see significant variability in the results that appear to be due to the varying environmental conditions. I see frames 5800-6800 have a low error variance which corresponds to a period in which the vehicle is stopped at a light.
@@
A plot of the trajectory generated by IMU dead reckoning, panoramic GPS and INS + GPS integration is shown in Figure 10. From the plot I can see a significant bias in the gyroscope and accelerometer sensor that causes a drift in the trajectory. The INS + GPS trajectory near the panoramic GPS trajectory but at instances show vast deviation from the road when plotted and analysed visually. There is a scope of improvement in the trajectory generation by filtering out erroneous IMU data, initial calibration of the INS system and better filter tuning based on sensor characteristics.
@@ VI Future Work
There are several aspects of my system that I believe may be significantly improved upon. One area which I did not focus on in this paper is implementing my pipeline directly on a mobile phone running in real-time, however many of the components of my pipeline are already present in many mobile devices. Visual-Inertial Odometry and Extended Kalman Filters are, for example, present on most iOS devices with ARKit and run in real-time. Feature extraction and matching has been demonstrated to run in real-time at 30fps on mobile phones [13] and newer implementations of SIFT and nearest-neighbor matching on the GPU has shown these both of these steps can be performed in less than 4ms combined [10]. Performing real-time local bundle adjustment is more challenging, however although implementations have been demonstrated to run in 500ms on a mobile phone [12] and 230ms on a GPU [11]. Finally, while I found VLDs to perform robust outlier rejection for my pipeline, I am unaware of any work to accelerate K-VLD filtering or perform it on a mobile device. Future work may examine the performance of K-VLD filtering on mobile devices and the potential to perform these computations on a remote server and transmit the final results to the client mobile device.
@@ VII Conclusion
my initial motivation for this project was that the sensor suite typically required for autonomous navigation is expensive and inaccessible for the vast majority of the population with motor vehicles. To address this issue, I built a system for accurate localization that is not gated by cost and would be easily accessible to anyone with a car. Using only a monocular camera, IMU, and GPS from a smartphone, along with panoramas retrieved from Google Street View, my method is able to compute a localization and trajectories with visual inertial odometry. Importantly, my method is able to localize with lane level accuracy, which is better than GPS alone, especially in dense city environments with large buildings or structures that obstruct RF signals. Additionally, with advancement in smartphone sensor technology and general smartphone availability, my method will only become more accurate and accessible.